# Application of Non-Linear Noise Regression in the Virgo Detector

### v3

@@ Abstract
Since the first detection of gravitational waves (GWs) in 2015, the International Gravitational-wave Network has made substantial strides in improving the sensitivity of ground-based detectors. Despite these advancements, many GW signals remain below the detection threshold due to environmental noise that limits sensitivity. In recent years, algorithms such as DeepClean have been developed to estimate and remove contamination from various noise sources, addressing linear, non-linear, and non-stationary coupling mechanisms. In this paper, we present noise reduction in the Virgo detector using DeepClean, serving as a preliminary step toward integrating Virgo into the online noise reduction pipeline for the O5 observing run. Our results demonstrate the applicability of DeepClean in Virgo O3b data, where noise was reconstructed from a total of 225 auxiliary witness channels. These channels were divided into 13 subsets, each corresponding to a specific frequency band, with training and subtraction performed layer-wise in a sequential manner. We observe that the subtraction improves the binary neutron star inspiral range by up to 1.3 Mpc, representing an approximately 2.5% increase. To ensure robust validation, we conduct an injection study with binary black hole waveforms. Matched-filter analyses of the injections showed an average improvement of 1.7% in the recovered SNR, while parameter estimation confirmed that DeepClean introduces no bias in the recovered parameters. The successful demonstration provides a pathway for online non-linear noise subtraction in Virgo in the future observing runs.
2D
two–dimensional
3D
three–dimensional
aLIGO
Advanced LIGO
AdVirgo
Advanced Virgo
ASD
amplitude spectral density
BAYESTAR
BAYESian TriAngulation and Rapid localization
BBH
Binary Black hole
BH
Black Hole
BNS
Binary Neutron star
CBC
Compact Binary Coalescence
CNN
Convolutional Neural Network
EM
electromagnetic
EOS
equation of state
FAR
False Alarm Rate
FOV
field of view
GCN
Gamma-ray Coordinates Network
GR
General Relativity
GRANDMA
Global Rapid Advanced Network Devoted to Multi-messenger Addicts
GRB
gamma-ray burst
GW
Gravitational-wave
gwemopt
Gravitational-wave Electromagnetic Optimization
GWTC-3
Gravitational Wave Transient Catalogue 3
$H_{0}$
Hubble-Lemaître constant
HEALPix
Hierarchical Equal Area isoLatitude Pixelization
IGWN
International Gravitational-wave Network
ISCO
innermost stable circular orbit
KAGRA
KAmioka GRAvitational–wave observatory
KDE
kernel density estimator
KN
kilonova
LHO
Laser Interferometer Observatory (LIGO) Hanford Observatory
LIGO
Laser Interferometer GW Observatory
LISA
The Laser Interferometer Space Antenna
LLO
LIGO Livingston Observatory
LRR
Living Reviews in Relativity
LSC
LIGO Scientific Collaboration
LSST
Large Synoptic Survey Telescope
$M_{\odot}$
sun mass
MLE
maximum likelihood (ML) estimator
ML
maximum likelihood
NIR
near infrared
NMMA
Nuclear-physics and Multi-Messenger Astrophysics
MSE
mean square error
NSBH
neutron star—black hole
NSBH
neutron starBlack Hole
NS
neutron star
O1
LIGO’s first observing run
O2
LIGO/Virgo’s second observing run
O3
LIGO/Virgo’s third observing run
O4
LIGO/Virgo/KAGRA’s fourth observing run
O5
LIGO/Virgo/KAGRA’s fifth observing run
O6
LIGO/Virgo/KAGRA/LIGO–India’s fifth observing run
PSD
power spectral density
PDB
Power Law + Dip + Break
PDB/GWTC-3
Power Law + Dip + Break/GWTC-3
SED
spectral energy distribution
SN
supernova
SN II
Type II supernova (SN)
SN Ia
Type Ia SN
SN IIb
Type IIb SN
SN Ib
Type Ib SN
SN Ic
Type Ic SN
SN Ic/Ib
Type Ic/b SN
SNR
signal–to–noise ratio
SVD
singular value decomposition
TOO
target–of–opportunity
ULTRSAT
Ultraviolet Transient Astronomy Satellite
UV
ultraviolet
UVEX
Ultraviolet Explorer
Roman
Nancy Grace Roman Space Telescope
Rubin Observatory
Vera C. Rubin Observatory
WD
white dwarf
ZTF
Zwicky Transient Facility
1 Introduction
Gravitational-wave (GW) events, such as GW150914 (Abbott et al., 2016), GW170817 (Abbott et al., 2017), GW190425 (Abbott et al., 2020b), and GW200105 (Abbott et al., 2020b), among others, have significantly improved our understanding of compact binary mergers (Abbott et al., 2021a, 2023). However, many potential GW signals remain undetected just below the noise floor, awaiting the successful mitigation of environmental noise sources. In the absence of a GW, the detectors may show fluctuations due to various factors that contribute to the noise of the system, including fundamental, technical, and environmental sources. These noise sources contribute with different and time-varying characteristics in amplitude and frequency. The characterization of the various noise components is crucial to understanding and improving the performance of a detector, in particular when identifying authentic GW signals (Acernese et al., 2023a; Buikema et al., 2020). Fundamental noise includes inherent limitations of detector design materials and quantum mechanics. Thermal noise in the mirror coatings is a fundamental noise in the frequency band of approximately 60-300 Hz. Environmental noise covers external influences like seismic noise, atmospheric noise, external thermal fluctuations, and Newtonian noise (Martynov et al., 2016). Technical noise comes from the design and operational aspects of the detector.
The various noise sources make it difficult to clearly discern GW signals, mitigating or subtracting noise within the detectors may expose these missed signals to detection algorithms. To potentially reveal new signals or otherwise improve the signal–to–noise ratio (SNR) of known ones, both the Laser Interferometer GW Observatory (Advanced LIGO; LSC et al. 2015) and Advanced Virgo (Acernese et al., 2014) record several thousands of witness channels (probes and sensors) (Allocca et al., 2020a; Acernese et al., 2023a). These witness channels independently are sensitive to noise from various sources. Some of them collect data to characterize environmental noise sources that couple with the GW readout channel and could therefore be used to subtract noise sources. Only non-fundamental noise can be subtracted. Thus, we prioritize mitigating measurable environmental and technical noise. Through the classification of noise into removable and non-removable categories, our objective is to enhance the sensitivity of GW detectors, ensuring a clearer distinction between real GW signals and noise artifacts.
In the past, some noise subtraction methods, such as Wiener filtering, widely used to subtract linearly coupled noise sources, can face challenges in the presence of non-stationary noise (Davis et al., 2019). Furthermore, non-linear noise, arising from systems in which the output is not directly proportional to the input, creates complex and unpredictable noise patterns. This non-linearity can be due to a multitude of factors, including complex physical processes within the detector or environmental influences (Driggers et al., 2019). While non-linear noise reflects a non-proportional relationship, non-stationary noise refers instead to time-dependent variations in an otherwise linear system (Vajente et al., 2020). On the other hand, non-stationary noise is characterized by its statistical properties that change over time, adding a layer of complexity to data analysis and signal processing. The dynamic variability of the noise makes it particularly challenging to model and subtract from the data.
At present, there are several narrow-band and broadband frequency ranges where the sensitivity is significantly made worse than that expected by the design. Low-frequency (typically around 10-100 Hz) sensitivity is particularly important for many reasons (Ormiston et al., 2020; Vajente et al., 2020). These include the possible detections of high-mass binary mergers at low frequencies, stochastic GW background detection which significantly benefits from the excess GW power in low-frequencies (Christensen, 2018; Abbott et al., 2021b), and the pre-merger detectability of Binary Neutron star (BNS) mergers (Abbott et al., 2017) as it allows us to anticipate potential electromagnetic (EM) counterparts (Petrov et al., 2022; Kiendrebeogo et al., 2023). Furthermore, many GW signals that are hidden in noise, would become detectable when noise is adequately minimized in any frequency band.
@@ 1 Introduction
DeepClean is a deep-learning infrastructure developed to denoise the GWs strain data using witness channels (Ormiston et al., 2020; Saleem et al., 2023). It has a 1D Convolutional Neural Network (CNN) architecture, with the time-series data from multiple witness sensors as inputs and a single noise prediction as the output. Due to its complex architecture, DeepClean is capable of subtracting non-linear and non-stationary noise from the GW data. The performance of DeepClean on LIGO detectors has already been demonstrated in detail in the aforementioned publications (Ormiston et al., 2020; Saleem et al., 2023).
@@
In this work, we focus on the Virgo detector, where some EM noise comes from electrical sources, reaching their peak at the 50 Hz power line frequency. This electrical noise induces modulations in the main GW signal, leading to the generation of symmetric sidebands around 50 Hz. During O3, a feed-forward loop process was used to remove the Virgo mains and its associated sidebands (Allocca et al., 2020b). We also aim to demonstrate the flexibility of the architecture in to be applicable on detectors with different designs. In addition, in this work, we also demonstrate a multi-layer multi-band training approach (simply known as multi-training) unlike the subtractions done in Saleem et al. (2023) which were done for a single frequency band with a single layer of training.
@@
In Sec. 2, we describe the methodology implemented in the DeepClean algorithm, focusing on noise estimation and reduction using witness channels. Sec. 3 gives a discussion of the application of DeepClean on Virgo’s O3b (the second part of the third observing run of LIGO-Virgo collaboration) data, including the training process and noise subtraction. The results, including analyses of individual frequency bands and the multi-training approach, are presented in Sec.4. Our conclusion and future prospects are summarized in Sec. 5.
2 Noise Subtraction in gravitational wave data
2.1 Popular subtraction techniques
Wiener filtering—This is a fundamental filtering technique that is effective when the noise is stationary and noise coupling is linear. It aims at minimizing the mean square error (MSE) between estimated and true signals (Wiener, 1949; Davis et al., 2019). Its robustness is evident in GW event analysis by the International Gravitational-wave Network (IGWN) (Abbott et al., 2020a). However, its performance becomes sub-optimal for non-stationary and non-linear noise driven by complex environmental and instrumental fluctuations. This highlights the need for adaptive noise reduction strategies in GW astronomy.
NonSENS—Non-Stationary Estimation of Noise Subtraction (NonSENS) employs an Infinite Impulse Response (IIR) filter to remove both stationary and non-stationary noise in GW detectors (Vajente et al., 2020). This method effectively addresses noise arising from slow interferometer movements, such as angular fluctuations, which pose challenges to traditional techniques. By utilizing witness signals to monitor noise and its modulation over time, NonSENS achieves stable, parametric subtraction of non-stationary noise. It leverages data science techniques, such as gradient-based optimization (Kingma & Ba, 2014), to estimate the coupling coefficients, thereby overcoming the complexities that hinder standard noise reduction methods. Additionally, the analytical filter, with its limited number of coefficients, avoids the computational challenges and interpretability issues often associated with deep neural networks.
Adaptive Feed-Forward—The feed-forward technique in Advanced Virgo is designed to mitigate 50 Hz noise by using a phase signal from the Uninterruptible Power Supply (UPS) as a witness channel. First, the system measures the 50 Hz noise through the UPS. Then, it dynamically adjusts the gain and phase to match the noise observed in the differential arm length (DARM) channel. A 50 Hz resonant filter is applied to isolate and target this specific frequency. Finally, the corrected signal is injected into DARM to cancel out the noise. This adaptive approach continuously adjusts to variations in the noise environment, ensuring efficient and reliable suppression of the 50 Hz disturbance.
2.2 DeepClean
@@
Neural networks have emerged as the leading approach for noise reduction due to their adaptability and robustness in managing complex data. These models effectively distinguish between signal and noise, thereby enhancing the overall performance of GW detectors. DeepClean uses a CNN architecture to measure the correlation between auxiliary sensors and GW data (Ormiston et al., 2020; Saleem et al., 2023), to mitigate environmental and instrumental noise off the GW data. Figure 1 summarizes the workflow. The architecture features a symmetric one-dimensional auto-encoder with multiple downsampling and upsampling layers, tailored to the sampling frequency and witness channel count. This configuration enables noise prediction across different data dimensions and uses batch normalization and tanh activation functions to boost generalization. The trainable weights of the neural network determine the form of the noise coupling. In the most general case, it can capture non-linear couplings by virtue of the non-linear activation functions included in the neural network. With the help of auxiliary sensors tracking the changes in coupling features over time, it can in principle also capture non-stationary noise couplings. The training minimizes a loss function based on the noise spectrum ratio of cleaned to original strain across all frequency bins.
@@
where $PSD_{h}[k]$ represents the power spectral density (PSD) derived from the GW strain data, and $PSD_{r}[k]$ represents the PSD of the residual signal, $r(t)$. The index $k$ runs over all the $N$ frequency bins, uniformly spaced between $f_{low}$ and $f_{high}$. To optimize the weights $(\vec{\theta})$, DeepClean employs the ADAM optimizer, which is an efficient algorithm to minimize the loss function in the gradient space (Kingma & Ba, 2014).
3 A Virgo O3b Mock Data Challenge
@@
We apply DeepClean on Virgo O3b data and assess the performance using various metrics. We examined a continuous stretch of GW strain data Hrec_hoft_raw_20000Hz and the useful witness channels, spanning two days, 18 hours and 15 minutes, from February 7, 2020, 16:19:27 UTC to February 10, 2020, 10:35:01 UTC. This period was selected for its extensive coverage and the significant coherence observed between strain and auxiliary channels in the Virgo O3b dataset which is essential for the effective deployment and evaluation of DeepClean. The models were trained in 1024 s, 2048 s, and 4096 s segments with the training performed once for every 100,000 s of data, after which the model’s performance in noise subtraction significantly decreased, necessitating retraining. Each model was used to clean the corresponding segment length. We first focus noise reduction specifically within three frequency bands: 98–110 Hz, 142–162 Hz, and 197–208 Hz. We target these bands due to their excessive noise, which is appropriate for testing the effectiveness of the algorithm and its ability in enhancing the recovery of astrophysical signals in the data. Noise reduction in these bands are particularly interesting in the context of pre-merger detections of Compact Binary Coalescence (CBC) events, especially binary neutron star mergers and other low mass black hole binaries.
3.1 Witness sensors
A comprehensive analysis of witness channels was first performed within the specific frequency bands.
@@ 3.1 Witness sensors
To determine which channels are the most important, we consider the coherence between witness channels and the GW strain channel. We consider the output of the so-called BruCo algorithm (a brute-force coherence computation tool among all channels in Virgo; Acernese et al. 2023b), which run automatically during the O3b, to select channels whose coherence exceeded 0.5 (50% of the maximum possible coherence). Additionally, we required that this coherence excess occurred at least five times during February and March 2020. This ensures that each witness channel shows sustained coherence levels across multiple instances. Each witness channel is tested for its noise removal effects and those deemed to not contribute are removed. Below we describe the main witness channels selected for each frequency band.
@@
142 to 162 Hz: Given the complex contributions of 167 witness channels in this frequency range to noise reduction, a full listing is omitted. Key acronyms like ASC (Alignment Sensing and Control), INJ (Injection), ACT (Actuation), and ENV (Environment) are emphasized to highlight the main systems that maintain the operating integrity of the interferometer (Acernese et al., 2023a). The complete list of witness channels can be found here  [^3] https://github.com/weizmannk/Virgo-DeepClean/blob/main/config/witnesses/single-layers/witnesses_142-162_Hz.ini.
3.2 Training and cleaning

### v4

@@ Abstract
This work presents the first demonstration of non-linear noise regression in the Virgo detector using deep learning techniques. We use DeepClean, a convolutional autoencoder previously shown to be effective in denoising LIGO data, as our tool for modeling and subtracting environmental and technical noise in Virgo. The method uses auxiliary witness channels to learn correlated noise features and remove them from the strain data. For this study, we apply DeepClean to Virgo O3b data, using 225 witness channels selected across 13 targeted frequency bands. Our analysis confirms the presence of non-linear couplings in the subtracted noise, highlighting the importance of DeepClean-like tools in capturing such effects. We observe up to a 1.3 Mpc improvement in the binary neutron star inspiral range ($\sim$2.5% gain), and an average increase of 1.7% in the recovered signal-to-noise ratio for injected binary black hole signals. Parameter estimation studies further confirm that DeepClean does not introduce bias in the recovery of source parameters. These results demonstrate the robustness of DeepClean on Virgo data and support its adoption in real-time noise subtraction frameworks for future observing runs.
preprint: APS/123-QED
I Introduction
Gravitational-wave (GW) events, such as GW150914 [1], GW170817 [2], GW190425 [3], and GW200105 [3], among others, have significantly improved our understanding of compact binary mergers [4, 5]. However, many potential GW signals remain undetected just below the noise floor, awaiting the successful mitigation of environmental noise sources.
In the absence of a GW, the detectors may show fluctuations due to various factors that contribute to the noise of the system, including fundamental, technical, and environmental sources. These noise sources contribute with different and time-varying characteristics in amplitude and frequency. The characterization of the various noise components is crucial to understanding and improving the performance of a detector, in particular when identifying authentic GW signals [6, 7]. Fundamental noise includes inherent limitations of detector design materials and quantum mechanics. Thermal noise in the mirror coatings is a fundamental noise in the frequency band of approximately 60-300 Hz. Environmental noise covers external influences like seismic noise, atmospheric noise, external thermal fluctuations, and Newtonian noise [8]. Technical noise comes from the design and operational aspects of the detector.
GW detectors are affected by various noise sources. Although the $h(t)$ readouts are similar, they differ in arm length (4 km for LIGO and 3 km for Virgo), electronic components, optical configuration, control design, and geographic location, which expose them to different operating conditions and to distinct noise environments that influence their sensitivities. In this work, we present the first application of DeepClean to Virgo data, a dedicated noise subtraction effort aimed at improving the detector sensitivity.
The presence of various noise sources make it difficult to clearly discern GW signals, mitigating or subtracting noise within the detectors may expose these missed signals to detection algorithms. To potentially reveal new signals or otherwise improve the signal-to-noise ratio (SNR) of known ones, both the Laser Interferometer GW Observatory (Advanced LIGO; [9]) and Advanced Virgo [10] record several thousands of witness channels (probes and sensors) [11, 6]. These witness channels independently are sensitive to noise from various sources. Some of them collect data to characterize environmental noise sources that couple with the GW readout channel and could therefore be used to subtract noise sources.
Only non-fundamental noise can be subtracted. Thus, we prioritize mitigating measurable environmental and technical noise. Through the classification of noise into removable and non-removable categories, our objective is to enhance the sensitivity of GW detectors, ensuring a clearer distinction between real GW signals and noise artifacts.
Noise subtraction methods such as Wiener filtering have been widely used to subtract linearly coupled noise sources, however they face limitations in the presence of non-linear and non-stationary noise [12].
At present, there are several narrow-band and broadband frequency ranges where the sensitivity is significantly made worse than that expected by the design. Low-frequency (typically around 10-100 Hz) sensitivity is particularly important for many reasons [13, 14]. These include the possible detections of high-mass binary mergers at low frequencies, stochastic GW background detection which significantly benefits from the excess GW power in low-frequencies [15, 16], and the pre-merger detectability of binary neutron star (BNS) mergers [2] as it allows us to anticipate potential electromagnetic (EM) counterparts [17, 18]. Furthermore, many GW signals that are hidden in noise, would become detectable when noise is adequately minimized in any frequency band.
@@ I Introduction
DeepClean is a deep-learning infrastructure developed to denoise the GWs strain data using witness channels [13, 19]. It has a 1D Convolutional Neural Network (CNN) architecture, with the time-series data from multiple witness sensors as inputs and a single noise prediction as the output. Due to its complex architecture, DeepClean is capable of modeling and subtracting non-linear noise couplings from the GW data. DeepClean is also a light-weight architecture which takes only a few minutes to train on roughly one hour of data and few seconds to minutes to perform cleaning on several hours of data (for example, on a standard NVIDIA A100 Tensor Core GPU), potentially making the analysis time much smaller than the data duration. This makes it feasible to retrain DeepClean so often that the non-stationary features can be efficiently captured. The performance of DeepClean on LIGO detectors has already been demonstrated in detail in the aforementioned publications [13, 19].
@@
In this work, we focus on the Virgo detector, where some EM noise comes from electrical sources, reaching their peak at the 50 Hz power line frequency. This electrical noise induces modulations in the main GW signal, leading to the generation of symmetric sidebands around 50 Hz. During O3, a feed-forward loop process was used to remove the Virgo mains and its associated sidebands [20]. We also aim to demonstrate the flexibility of the architecture in to be applicable on detectors with different designs. In addition, in this work, we also demonstrate a multi-layer multi-band training approach (simply known as multi-training) unlike the subtractions done in [19] which were done for a single frequency band with a single layer of training.
@@
In Sec. II, we describe the methodology implemented in the DeepClean algorithm, focusing on noise estimation and reduction using witness channels. Sec. III gives a discussion of the application of DeepClean on Virgo’s O3b (the second part of the third observing run of LIGO-Virgo collaboration) data, including the training process and noise subtraction. The results, including analyses of individual frequency bands and the multi-training approach, are presented in Sec.III.3. Our conclusion and future prospects are summarized in Sec. V.
II DeepClean as a tool for non-linear noise regression
Several noise regression techniques have been developed to mitigate environmental and instrumental disturbances in GW detectors. These include classical linear methods such as Wiener filtering [21, 12, 22], and adaptive feed-forward control schemes [20], as well as more recent ML-supported parametric approaches like NonSENS [14]. Beyond these, a growing body of work explores ML-based denoising strategies [23, 24, 25, 26, 27], positioning DeepClean within the evolving landscape of data-driven noise regression. Moreover, ML is also increasingly applied to the identification of electromagnetic counterparts of GW sources [28, 29, 30].
See Appendix .1 for a short summary of some of the popular methods. While each method has its advantages, recent developments have shown that deep learning tools offer a powerful framework for capturing complex, non-linear, and potentially non-stationary noise couplings. In this study, we use DeepClean, a convolutional autoencoder previously demonstrated in LIGO data [13, 19], to perform non-linear noise regression in the Virgo detector. We describe the architecture and training methodology of DeepClean in detail below.
@@ II DeepClean as a tool for non-linear noise regression
DeepClean uses a CNN architecture to measure the correlation between different auxiliary sensors that is coupled to the GW data [13, 19], to mitigate environmental and instrumental noise off the GW data. Figure 1 summarizes the workflow. The architecture features a symmetric one-dimensional auto-encoder with multiple downsampling and upsampling layers, tailored to the sampling frequency and witness channel count. This configuration enables noise prediction across different data dimensions and uses batch normalization and tanh activation functions to boost generalization. The trainable weights of the neural network determine the form of the noise coupling. In the most general case, it can capture non-linear couplings by virtue of the non-linear activation functions included in the neural network [13]. With the help of auxiliary sensors tracking the changes in coupling features over time, it can in principle also capture non-stationary noise couplings. The training minimizes a loss function based on the noise spectrum ratio of cleaned to original strain across all frequency bins.
@@
where $PSD_{h}[k]$ represents the power spectral density (PSD) derived from the original strain data, and $PSD_{r}[k]$ represents the PSD of the residual strain, $r(t)$. The index $k$ runs over all the $N$ frequency bins, uniformly spaced between $f_{low}$ and $f_{high}$. To optimize the weights $(\vec{\theta})$, DeepClean employs the ADAM optimizer[31].
III A Virgo O3b Mock Data Challenge
We apply DeepClean on Virgo O3b data and assess the performance using various metrics. We examined a continuous stretch of GW strain data Hrec_hoft_raw_20000Hz and the useful witness channels, spanning two days, 18 hours and 15 minutes, from February 7, 2020, 16:19:27 UTC to February 10, 2020, 10:35:01 UTC. This period was selected for its extensive coverage and the significant coherence observed between strain and auxiliary channels in the Virgo O3b dataset which is essential for the effective deployment and evaluation of DeepClean.
The models were trained on segments of 1024 s, 2048 s, and 4096 s durations, with the segment length empirically chosen to balance computational cost with training stability. In practice, longer training segments tend to yield more robust learning by emphasizing persistent features of the noise couplings rather than short-term variations. To account for evolving coupling behavior and auxiliary channel reliability, the models were retrained approximately once every $10^{5}$ s of data. Beyond this cadence, subtraction performance degraded noticeably, necessitating periodic retraining.
@@ III A Virgo O3b Mock Data Challenge
Each model was used to clean the corresponding segment length. We first focus noise reduction specifically within three frequency bands: 98–110 Hz, 142–162 Hz, and 197–208 Hz. We target these bands due to their excessive noise, which is appropriate for testing the effectiveness of the algorithm and its ability in enhancing the recovery of astrophysical signals in the data. Noise reduction in these bands are particularly interesting in the context of pre-merger detections of compact binary coalescence (CBC) events, especially binary neutron star mergers and other low mass black hole binaries.
III.1 Witness sensors
@@
A comprehensive analysis of witness channels was first performed within the specific frequency bands. To determine which channels are the most important, we consider the coherence between witness channels and the GW strain channel. We consider the output of the so-called BruCo algorithm (a brute-force coherence computation tool among all channels in Virgo; [32]), which run automatically during the O3b, to select channels whose coherence exceeded 0.5 (50% of the maximum possible coherence). Additionally, we required that this coherence excess occurred at least five times during February and March 2020. This ensures that each witness channel shows sustained coherence levels across multiple instances. Each witness channel is tested for its noise removal effects and those deemed to not contribute are removed. The choice of a $0.5$ threshold represents a trade-off: lowering the threshold would include many weakly correlated channels, which in our tests did not improve the subtraction performance but significantly increased the training complexity and computational cost. Conversely, setting the threshold too high risks excluding moderately coupled channels that still provide useful noise regression. The adopted criterion therefore balances computational efficiency with denoising performance, ensuring a manageable set of channels with consistent contribution. Below we describe the main witness channels selected for each frequency band.
@@
142 to 162 Hz: Given the complex contributions of 167 witness channels in this frequency range to noise reduction, a full listing is omitted. Key acronyms like ASC (Alignment Sensing and Control), INJ (Injection), ACT (Actuation), and ENV (Environment) are emphasized to highlight the main systems that maintain the operating integrity of the interferometer [6]. The complete list of witness channels can be found here  [^3] https://github.com/weizmannk/Virgo-DeepClean/blob/main/config/witnesses/single-layers/witnesses_142-162_Hz.ini.
III.2 Training and cleaning